# Constrained Denoising, Empirical Bayes, and Optimal Transport

### v1

@@ 4 Empirical Constrained Bayes Denoising
Specifically, we suppose that $\delta_{\ast}$ is one of $\delta_{\mathcal{VCB}},\delta_{\mathcal{DCB}}$, or $\delta_{\mathcal{GCB}}$, and we aim to develop an EB estimator $\hat{\delta}_{\ast}$ of $\delta_{\ast}$. All of our estimators arise by applying some further processing to a denoiser $\hat{\delta}_{\mathcal{B}}$ which is already known to be a good approximation of the Bayes optimal (that is, unconstrained) denoiser $\delta_{\mathcal{B}}$. More precisely, if we assume a priori that $\hat{\delta}_{\mathcal{B}}$ satisfies
$\frac{1}{n}\sum_{i=1}^{n}\left\|\hat{\delta}_{\mathcal{B}}(Z_{i})-\delta_{\mathcal{B}}(Z_{i})\right\|^{2}\overset{\mathbb{P}}{\to}0,$ (4.1)
then we can construct $\hat{\delta}_{\ast}$ from $\hat{\delta}_{\mathcal{B}}$ in such a way that we have $\hat{\delta}_{\ast}\to\delta_{\ast}$ as $n\to\infty$; moreover the rate of convergence of $\hat{\delta}_{\ast}$ to $\delta_{\ast}$ depends explicitly on the rate of convergence of $\hat{\delta}_{\mathcal{B}}$ to $\delta_{\mathcal{B}}$ given in (4.1) above. This flexibility in choosing $\hat{\delta}_{\mathcal{B}}$ is important for the broad applicability of the methodology; the reader can keep in mind the following examples for how EB denoisers may arise. Here we follow Efron (2014) and distinguish between $G$-modeling and $F$-modeling approaches.
@@
Example 4.3 ($F$-modeling).
If the Bayes denoiser $\delta_{\mathcal{B}}$ depends only on the marginal distribution $F$ of $Z_{1},\ldots,Z_{n}$, then an EB approximation $\hat{\delta}_{\mathcal{B}}$ of $\delta_{\mathcal{B}}$ arises by modeling $F$ rather than $G$. This is possible when $P$ is a continuous exponential family with natural parameterization, due to Eddington’s/Tweedie’s formula (Dyson, 1926; Efron, 2011), but it is also possible in many special cases where similar formulas are available (Robbins, 1956, 1982; Cressie, 1982).
Example 4.4 (Exponential Family, Conjugate Prior).
Suppose in $d=m=1$ that $\{P_{\theta}\}_{\theta}$ is a mean-parameterized exponential family and that $G$ is a conjugate prior for $\{P_{\theta}\}_{\theta}$. We also assume that $\{P_{\theta}\}_{\theta}$ has a quadratic variance function, meaning there exists a quadratic polynomial $V:\mathbb{R}\to\mathbb{R}$ such that $P_{\theta}$ has variance $V(\theta)$ for all $\theta$; see Morris (1982) for more on such models. It is well-known (see Diaconis and Ylvisaker (1979)) that the Bayes denoiser $\delta_{\mathcal{B}}$ is a linear function in this setting; we have
where $\hat{\mu}:=\frac{1}{n}\sum_{i=1}Z_{i}$ and $\hat{s}^{2}:=\frac{1}{n-1}\sum_{i=1}^{n}(Z_{i}-\hat{\mu})^{2}$ are the empirical mean and variance of $Z_{1},\ldots,Z_{n}$. One can easily show in this setting that the convergence of $\hat{\delta}_{\mathcal{B}}$ to $\delta_{\mathcal{B}}$ has the parametric rate $n^{-1}$. For convenience, we give in Appendix B some explicit formulas for the denoiser above, in the usual parametric examples of interest (Gaussian, Poisson, Exponential, Geometric, etc.)
We address the case of variance constraints in Subsection 4.1, the case of distributional constraints in Subsection 4.2, and the case of general constraints in Subsection 4.3. The proofs of all results in this section can be found in Appendix E.
@@ 4.1 Variance-Constrained Denoising
Last, we assume that we are equipped with some EB approximation $\hat{\delta}_{\mathcal{B}}$ of $\delta_{\mathcal{B}}$, and that it satisfies the following quantitative form of (4.1): For some specified sequence $\alpha_{n}\to 0$, we have
@@
This leads us to the main result, whose proof appears in Appendix E.1:
@@
$\frac{1}{n}\sum_{i=1}^{n}\left\|\hat{\delta}_{\mathcal{VCB}}(Z_{i})-\delta_{\mathcal{VCB}}(Z_{i})\right\|_{2}^{2}=O_{\mathbb{P}}\left(\alpha_{n}\vee n^{-2(q-2)/q}\right)$
as $n\to\infty$.
The conclusion of Theorem 4.5 for the denoiser given in Algorithm 1 is that the rate of convergence of variance-constrained Bayes denoising is determined by whichever of unconstrained Bayes denoising (line 4) and covariance estimation (line 8) is slower. Moreover, Appendix C contains a detailed discussion of the rates of convergence in the Gaussian and Poisson models arising from this result, and we observe that it is possible for either of these terms to dominate. In the case of a Gaussian likelihood and a subexponential latent variable distribution, we note that the resulting rate is nearly parametric, i.e. $n^{-1}$ up to logarithmic factors. In the case of an an exponential family likelihood and a conjugate prior, the resulting rate is exactly parametric.
@@ 4.2 Distribution-Constrained Denoising
From the characterization of $\delta_{\mathcal{DCB}}$ as the solution to the OT problem (3.5), it is clear that, in addition to the existence of an EB denoiser $\hat{\delta}_{\mathcal{B}}$, we also need to be able to estimate the distribution $F$, the distribution $G$, and the cost function $c_{G}$. (Recall its definition in (3.4).) We may estimate $F$ via the empirical distribution of $Z_{1},\ldots,Z_{n}$, $G$ as $\hat{G}_{n}$ via $G$-modeling (as in, e.g., Examples 4.1 and 4.2), and $c_{G}$ as a function of the given EB denoiser $\hat{\delta}_{\mathcal{B}}$; then we use these approximations in an OT problem paralleling (3.5), and the barycentric projection of its optimal coupling is an approximation of $\delta_{\mathcal{DCB}}$. (Recall from Section 2 that the barycentric projection recovers the OT map if the optimal coupling is supported on the graph of a function, and that it provides an approximation of the OT map otherwise.) We define this procedure explicitly in Algorithm 2, and we prove a rate of convergence as our main result below.
@@
$\frac{1}{n}\sum_{i=1}^{n}\left\|\hat{\delta}_{\mathcal{DCB}}(Z_{i})-\delta_{\mathcal{DCB}}(Z_{i})\right\|^{2}=O_{\mathbb{P}}\left(\alpha_{n}^{\nicefrac{{1}}{{2}}}\vee\beta_{n}\vee\gamma_{n}\right)$ (4.4)
as $n\to\infty$.
The conclusion of Theorem 4.6 for the denoiser given in Algorithm 2 is that the rate of convergence of distribution-constrained Bayes denoising is determined by whichever of unconstrained Bayes denoising (line 4), deconvolution (line 5), and convergence of the empirical Wasserstein distance is slowest. While in Appendix C.2 we give a detailed discussion of the resulting rates of convergence in some concrete settings of interest, we emphasize that $\beta_{n}$ typically dominates in nonparametric problems and that $\gamma_{n}$ typically dominates in parametric problems. The term $\alpha_{n}^{\nicefrac{{1}}{{2}}}$ is likely an artifact of the proof and we expect that it can be improved to $\alpha_{n}$ with further work, but we do not pursue this since it does not affect the rates in cases of interest.
@@ 4.3 General-Constrained Denoising
Lastly, we study problem ($\mathcal{GCB}$) on denoising subject to general constraints. Since in Theorem 3.12 we provided sufficient conditions for the existence of a unique solution to problem ($\mathcal{GCB}$), we presently develop an EB methodology for targeting this oracle denoiser. Methodologically, we regard ($\mathcal{GCB}$) as a way to tradeoff between problems ($\mathcal{VCB}$) and ($\mathcal{DCB}$); the former only encodes limited information about the latent variable distribution (the first two moments), the latter suffers from slow rates of convergence (due to the difficulty of the deconvolution problem) and may have larger denoising risk. Often some intermediate behavior may be desired.
We describe our proposed procedure in Algorithm 3, which is similar to the approach we previously used for the distribution-constrained problem ($\mathcal{DCB}$). Since we can estimate the marginal distribution $F$ via $\bar{F}_{n}:=\frac{1}{n}\sum_{i=1}^{n}\delta_{Z_{i}}$ and the cost function $c_{G}$ via $\hat{c}_{n}$ through $\hat{\delta}_{\mathcal{B}}$, it only remains to approximate the constraints of the linear program (3.8). Assuming that this is possible (that is, that the integrals of $\Psi$ with $\hat{G}_{n}$ converge to the integrals of $\Psi$ with $G$), we may let $\hat{\delta}_{\mathcal{GCB}}$ denote the barycentric projection of an optimal solution to the corresponding linear program. Note also that we restrict the $\eta$ variable of the linear program to lie in a compact set $K\subseteq\mathbb{R}^{m}$ (line 6).
In order to give a rate of convergence of $\hat{\delta}_{\mathcal{GCB}}$ to $\delta_{\mathcal{GCB}}$, we need to present a set of assumptions. To set this up, we first assume (ZC), (BC), (C), and (QG) so that Theorem 3.12 is in effect; then, we recall that there is a well-defined projected latent variable distribution which we denote $H$. Second, we assume
in probability as $n\to\infty$. For $\psi_{\ell}$ with at most quadratic growth, such convergence follows from (EC) because of Villani (2009, Definition 6.8). For $\psi_{\ell}$ whose growth is polynomial of degree $p\geq 2$, it is sufficient to additionally require $W_{p}(\hat{G}_{n},G)=o_{\mathbb{P}}(1)$ on top of (EC).
$W_{2}^{2}(\hat{H}_{n},H)=O_{\mathbb{P}}(\delta_{n}).$ (D)
which is analogous to (EW).
5:     $\hat{G}_{n}\leftarrow$ estimate of $G$
8:                 over         probability measures $\pi\in\mathcal{P}(\mathbb{R}^{d}\times K)$
9:                 with        $\pi(\{Z_{i}\}\times K)=\frac{1}{n}$ for all $1\leq i\leq n$
12:     return $\hat{\delta}_{\mathcal{GCB}}$
13:end procedure
Finally, these considerations lead us to the following result.
Theorem 4.7.
$\frac{1}{n}\sum_{i=1}^{n}\left\|\hat{\delta}_{\mathcal{GCB}}(Z_{i})-\delta_{\mathcal{GCB}}(Z_{i})\right\|^{2}=O_{\mathbb{P}}(\alpha_{n}^{\nicefrac{{1}}{{2}}}\vee\delta_{n}\vee\varepsilon_{n})$ (4.5)
as $n\to\infty$.
The conclusion of Theorem 4.7 for the denoiser $\hat{\delta}_{\mathcal{GCB}}$ given in Algorithm 3 is that the rate of convergence of general-constrained Bayes denoising is determined by whichever of unconstrained Bayes denoising), convergence of the projected latent variable distributions, and convergence of the empirical Wasserstein distance is slowest. The rate of convergence of the projected latent variable distributions, described by $\delta_{n}$, exactly quantifies the statistical tradeoff of interpolating between problems ($\mathcal{VCB}$) and ($\mathcal{DCB}$). In the cases we consider in Example 3.13, we may anticipate $\delta_{n}$ to decay at nearly parametric rates when $k=0$ (no constraints), or when $k=2$ (variance constraints) and the empirical Bayes denoiser may be estimated at fast rates (as discussed in Example 4.1), or inversely polylogarithmic in $n$ when $k\to\infty$ (and we effectively need to deconvolve the latent variable distribution).
@@ 5 Heterogeneous (Empirical) Constrained Bayes Denoising
Our results and methods so far have been developed under the assumption that all $n$ observations follow model (1.1). While this modeling setup covers many important applications, it excludes settings with heterogeneity in the likelihood, e.g., settings with varying precision in the observations. As one concrete example, suppose that conditional on $\Theta_{i}$, $Z_{i}$ is a sample average of $K_{i}$ i.i.d. observations $Z_{ij}$ with $Z_{ij}\sim\mathcal{N}(\Theta_{i},\sigma^{2})$. Then $Z_{i}\sim\mathcal{N}(\Theta_{i},\sigma_{i}^{2})$ with $\sigma_{i}^{2}=\sigma^{2}/K_{i}$. This setting is covered by (1.1) only if $\sigma_{i}^{2}$ does not depend on $i$, i.e., only when the sample sizes $K_{i}$ for all $i$ are identical. Yet, it is common in applications to encounter datasets with varying sample sizes.
To accommodate settings with heterogeneity (e.g., with heteroskedastic Gaussian observations above), we extend model (1.1) by including a heterogeneity parameter $\Xi\in\mathbb{R}^{p}$ in addition to the latent variable $\Theta$ and the observation $Z$:
@@
Here, $\{G_{\xi}\}_{\xi}$ is an unknown family of probability measures on $\mathbb{R}^{m}$ and $\{P_{\theta,\xi}\}_{\theta,\xi}$ is a known family of probability distributions on $\mathbb{R}^{d}$. Usually we will assume that $\{G_{\xi}\}_{\xi}$ does not depend on $\xi$, i.e., there is some unknown $G$ satisfying $G_{\xi}=G$ for all $\xi$. We assume that $\Theta$ is unobserved while $(Z,\Xi)$ is observed. We also consider the EB setting where we have i.i.d. samples $(\Theta_{1},\Xi_{1},Z_{1}),\ldots,(\Theta_{n},\Xi_{n},Z_{n})$ from (5.1) and we observe $(\Xi_{1},Z_{1}),\ldots,(\Xi_{n},Z_{n})$.
The reader should keep in mind the following two special cases of this setting:
  • (Heteroskedastic Gaussian.) For $m=d$, assume that the parameter $\theta$ is a vector in $\mathbb{R}^{m}$ and the parameter $\xi$ is a positive semi-definite $m\times m$ matrix, which we choose to denote by $\Sigma$. Let $P_{\theta,\Sigma}=\mathcal{N}(\theta,\Sigma)$. Note that this setting encompasses the situation with varying sample size (described above) by taking $m=d=1$ and $\Xi_{i}=\sigma^{2}/K_{i}$.
  • (Poisson with heterogeneous exposure.) For $m=d=1$, the parameter $\theta$ is a non-negative real number and the parameter $\xi$ is a non-negative real number, which we choose to denote by $\lambda$. Then let $P_{\theta,\lambda}=\textnormal{Poi}(\lambda\theta)$. We call $\lambda$ the exposure (Norberg, 1979). The intuition here is that we are interested in the mean of the $i$-th observation normalized by the exposure time, that is, $\Theta_{i}=\mathbb{E}[Z_{i}/\lambda_{i}]$; see Subsection 6.2 for an application to baseball in which $\lambda_{i}$ is the number of games played by the $i$-th player.
The unconstrained denoising problem in model (5.1) is as follows:
$\begin{cases}\textnormal{minimize}&\mathbb{E}\left[\|\delta(Z;\Xi)-\Theta\|^{2}\right]\\ \textnormal{over}&\delta:\mathbb{R}^{d}\times\mathbb{R}^{p}\to\mathbb{R}^{m}.\end{cases}$ (5.2)
$\delta_{\mathcal{B}}(z;\xi)=\mathbb{E}[\Theta\,|Z=z,\Xi=\xi].$
In the next subsections, we consider adding constraints to problem (5.2). While we will not prove rigorous results regarding asymptotic theory, we extend the strategies of the previous sections to accommodate the heterogeneity.
An important feature of heterogeneous constrained denoising is that we now must distinguish between two kinds of constraints: the so-called marginal constraints on the distribution of $\delta_{\mathcal{B}}(Z;\Xi)$, and the so-called conditional constraints on the distribution of $\delta_{\mathcal{B}}(Z;\Xi)$ given $\{\Xi=\xi\}$. We will consider problems of both forms for our later applications.
5.1 Variance-Constrained Denoising
5.1.1 Marginal Variance-Constrained Denoising
For the first approach, we require that the marginal mean and variance of the denoised data match the marginal mean and variance of the latent variable. That is, we solve problem (5.2) subject to the constraints
$\mathbb{E}[\delta(Z;\Xi)]=\mathbb{E}[\Theta]\qquad\textnormal{ and }\qquad\textnormal{Cov}(\delta(Z;\Xi))=\textnormal{Cov}(\Theta).$ (5.3)
This approach was also followed by Ghosh (1992) in a small area application with heterogeneous sample sizes across areas. By noticing that this is closely related to the setting of Section 3.1 but with observation $\tilde{Z}:=(Z,\Xi)$, it can be shown that the oracle denoiser is
$\hat{\mu}:=\int_{\mathbb{R}^{m}}\theta\textnormal{d}\hat{G}_{n}(\theta)\qquad\text{ and}\qquad\,\hat{A}:=\int_{\mathbb{R}^{m}}(\theta-\hat{\mu})(\theta-\hat{\mu})^{\top}\textnormal{d}\hat{G}_{n}(\theta).$ (5.4)
5.1.2 Conditional Variance-Constrained Denoising
Another possible constraint is that the conditional mean and covariance of $\delta(Z;\Xi)$ match the conditional mean and covariance of $\Theta$ given $\Xi$. In other words, we may solve problem (5.2) subject to the constraints
$\mathbb{E}[\delta(Z;\Xi)\,|\,\Xi]=\mathbb{E}[\Theta\,|\,\Xi]\qquad\textnormal{and }\qquad\textnormal{Cov}(\delta(Z;\Xi)\,|\,\Xi)=\textnormal{Cov}(\Theta\,|\,\Xi)\qquad\textnormal{ almost surely.}$. (5.5)
In this setting, it can be shown that the oracle denoiser is
where $\mu=\mathbb{E}[\Theta]$. Our EB approach will target the denoiser $\delta_{\mathcal{C}\mathcal{VCB}}$ by plugging in suitable estimates for $\mathbb{E}[\Theta],\textnormal{Cov}(\Theta\,|\,\Xi=\xi)$, and $\textnormal{Cov}(\delta_{\mathcal{B}}(Z,\Xi)\,|\,\Xi=\xi)$. If we additionally assume $\Theta\perp\!\!\!\!\perp\Xi$ (as we will assume in our later applications), then we may estimate $\mathbb{E}[\Theta]$ and $\textnormal{Cov}(\Theta)\equiv\textnormal{Cov}(\Theta\,|\,\Xi)$ in the same way as above, via the corresponding moments of the NPMLE given in (5.4). However, $\textnormal{Cov}(\delta_{\mathcal{B}}(Z,\Xi)\,|\,\Xi=\xi)$ must be handled more carefully. If we are in the setting of $G$-modeling and the measure $\hat{G}_{n}$ is finitely supported (which is true for the NPMLE), then we may approximate
via
$\sum_{j=1}^{r}w_{j}\int_{\mathbb{R}^{d}}(\hat{\delta}_{\mathcal{B}}(z,\xi)-\mu)(\hat{\delta}_{\mathcal{B}}(z,\xi)-\mu)^{\top}P_{\theta_{j},\xi}(\textnormal{d}z).$ (5.6)
5.1.3 Numerical Illustration
In Figure 5.1 we show the result of applying three denoisers to this data set. The marginal variance-constrained denoiser (third column) leads to the incorrect variance when applied only to the low-variance observations (second row) or the high-variance observations (third row). Contrarily, the conditional variance-constrained denoiser (fourth column) obtains the correct variance for both subsets of the data.
5.2 Marginal Distribution-Constrained Denoising
Next we seek to extend the distribution-constrained problem ($\mathcal{DCB}$) to the heterogeneous setting. To state the problem precisely, we aim to solve problem (5.2) subject to the constraint that the marginal distribution of the denoised data matches the marginal distribution of the latent variable, meaning
$\delta(Z;\Xi)\overset{\mathcal{D}}{=}\Theta.$ (5.7)
and we replace barycentric projection (line 12) by
Figure 5.1: Denoising in the heterogeneous setting of Section 5 to a toy data set. By rows, we show histograms of all of the observations (first row), the subset of low-variance observations (second row), and the subset of high-variance observations (third row). By columns, we show the raw data (first column), the EB denoised data set (second column), the marginal variance-constrained EB denoised data set (third column), and the conditional variance-constrained EB denoised data set (fourth column). In each plot we show the density of the latent variable distribution (blue) for convenience.
5.3 Marginal General-Constrained Denoising
$\mathbb{E}[\psi_{\ell}(\delta(Z,\Xi))]=\mathbb{E}[\psi_{\ell}(\Theta)]\quad\textnormal{ for all }1\leq\ell\leq k,$ (5.8)
where $\psi_{1},\ldots,\psi_{k}$ are some fixed functions. As in Subsection 3.3, the oracle denoiser in this problem is the composition of the Bayes denoiser $\delta_{\mathcal{B}}$ with an OT map from the distribution of $\delta_{\mathcal{B}}(Z,\Xi)$ to some distribution $H$ (whose integrals with respect to $\Psi$ agree with those of $G$) but which is not a priori known. Nonetheless, this denoiser can be approximated via applying the same modifications of Subsection 5.2 to Algorithm 3.

### v2

@@ 4 Empirical Constrained Bayes Denoising
Specifically, we suppose that $\delta_{\ast}$ is one of $\delta_{\mathcal{VCB}},\delta_{\mathcal{DCB}}$, or $\delta_{\mathcal{GCB}}$, and we aim to develop an EB estimator $\hat{\delta}_{\ast}$ of $\delta_{\ast}$. All of our estimators arise by applying some further processing to a denoiser $\hat{\delta}_{\mathcal{B}}$ which is already known to be a good approximation of the Bayes optimal (that is, unconstrained) denoiser $\delta_{\mathcal{B}}$. More precisely, if we assume a priori that $\hat{\delta}_{\mathcal{B}}$ satisfies $\frac{1}{n}\sum_{i=1}^{n}\left\|\hat{\delta}_{\mathcal{B}}(Z_{i})-\delta_{\mathcal{B}}(Z_{i})\right\|^{2}\overset{\mathbb{P}}{\to}0$, then we can construct $\hat{\delta}_{\ast}$ from $\hat{\delta}_{\mathcal{B}}$ in such a way that we have $\hat{\delta}_{\ast}\to\delta_{\ast}$ as $n\to\infty$; moreover the rate of convergence of $\hat{\delta}_{\ast}$ to $\delta_{\ast}$ depends explicitly on the rate of convergence of $\hat{\delta}_{\mathcal{B}}$ to $\delta_{\mathcal{B}}$. This flexibility in choosing $\hat{\delta}_{\mathcal{B}}$ is important for the broad applicability of the methodology; the reader can keep in mind the following examples for how EB denoisers may arise.
@@
In addition to $G$-modeling and smooth $G$-modeling, we note that other EB approaches are possible, including $F$-modeling and conjugate parametric models. In Appendix A we describe these in more detail, including some explicit formulas in the case of conjugate parametric models. Our results in this section are divided into variance constraints (Subsection 4.1), distributional constraints (Subsection 4.2), and general constraints (Subsection 4.3). The proofs of all results in this section can be found in Appendix G.
@@ 4.1 Variance-Constrained Denoising
Last, we assume that we are equipped with some EB approximation $\hat{\delta}_{\mathcal{B}}$ of $\delta_{\mathcal{B}}$, and that it satisfies the following quantitative notion of convergence: For some specified sequence $\alpha_{n}\to 0$, we have
@@
This leads us to the main result, whose proof appears in Appendix G.1:
@@
$\frac{1}{n}\sum_{i=1}^{n}\left\|\hat{\delta}_{\mathcal{VCB}}(Z_{i})-\delta_{\mathcal{VCB}}(Z_{i})\right\|_{2}^{2}=O_{\mathbb{P}}\left(\alpha_{n}\vee n^{-2(q-2)/q}\right)\qquad\textnormal{as}\qquad n\to\infty.$
The conclusion of Theorem 4.3 for the denoiser given in Algorithm 1 is that the rate of convergence of variance-constrained Bayes denoising is determined by whichever of unconstrained Bayes denoising (line 4) and covariance estimation (line 8) is slower. Moreover, Appendix B contains a detailed discussion of the rates of convergence in the Gaussian and Poisson models arising from this result, and we observe that it is possible for either of these terms to dominate. In the case of a Gaussian likelihood and a subexponential latent variable distribution, we note that the resulting rate is nearly parametric, i.e. $n^{-1}$ up to logarithmic factors. In the case of an an exponential family likelihood and a conjugate prior, the resulting rate is exactly parametric.
@@ 4.2 Distribution-Constrained Denoising
From the characterization of $\delta_{\mathcal{DCB}}$ as the solution to the OT problem (3.4), it is clear that, in addition to the existence of an EB denoiser $\hat{\delta}_{\mathcal{B}}$, we also need to be able to estimate the distribution $F$, the distribution $G$, and the cost function $c_{G}$. (Recall its definition in (3.3).) We may estimate $F$ via the empirical distribution of $Z_{1},\ldots,Z_{n}$, $G$ as $\hat{G}_{n}$ via $G$-modeling (as in, e.g., Examples 4.1 and 4.2), and $c_{G}$ as a function of the given EB denoiser $\hat{\delta}_{\mathcal{B}}$; then we use these approximations in an OT problem paralleling (3.4), and the barycentric projection of its optimal coupling is an approximation of $\delta_{\mathcal{DCB}}$. (Recall from Section 2 that the barycentric projection recovers the OT map if the optimal coupling is supported on the graph of a function, and that it provides an approximation of the OT map otherwise.) We define this procedure explicitly in Algorithm 2, and we prove a rate of convergence as our main result below. Details for implementing Algorithm 2 in practice can be found in Appendix E.
@@
$\frac{1}{n}\sum_{i=1}^{n}\left\|\hat{\delta}_{\mathcal{DCB}}(Z_{i})-\delta_{\mathcal{DCB}}(Z_{i})\right\|^{2}=O_{\mathbb{P}}\left(\alpha_{n}^{\nicefrac{{1}}{{2}}}\vee\beta_{n}\vee\gamma_{n}\right)\qquad\textnormal{as}\qquad n\to\infty.$ (4.3)
The conclusion of Theorem 4.4 for the denoiser given in Algorithm 2 is that the rate of convergence of distribution-constrained Bayes denoising is determined by whichever of unconstrained Bayes denoising (line 4), deconvolution (line 5), and convergence of the empirical Wasserstein distance is slowest. While in Appendix B.2 we give a detailed discussion of the resulting rates of convergence in some concrete settings of interest, we emphasize that $\beta_{n}$ typically dominates in nonparametric problems and that $\gamma_{n}$ typically dominates in parametric problems. The term $\alpha_{n}^{\nicefrac{{1}}{{2}}}$ is likely an artifact of the proof and we expect that it can be improved to $\alpha_{n}$ with further work, but we do not pursue this since it does not affect the rates in cases of interest.
@@ 4.3 General-Constrained Denoising
In Appendix C.2, we develop an EB methodology for targeting the oracle general-constrained denoiser $\delta_{\mathcal{GCB}}$. Methodologically, we regard ($\mathcal{GCB}$) as a way to tradeoff between problems ($\mathcal{VCB}$) and ($\mathcal{DCB}$); the former only encodes limited information about the latent variable distribution (the first two moments), the latter suffers from slow rates of convergence (due to the difficulty of the deconvolution problem) and may have larger denoising risk. Often some intermediate behavior may be desired. The rate of convergence of Theorem C.5 reflects the fact that ($\mathcal{GCB}$) can interpolate between ($\mathcal{VCB}$) and ($\mathcal{DCB}$).
@@ 5 Heterogeneous (Empirical) Constrained Bayes Denoising
Our results and methods so far have been developed under the assumption that all $n$ observations follow model (1.1). While this setting covers many important applications, it excludes settings with heterogeneity in the likelihood, e.g., settings with varying precision in the observations, as well as settings with heterogeneity in the prior. To accommodate such heterogeneity, we extend model (1.1) by including a heterogeneity parameter $\Xi\in\mathbb{R}^{p}$ in addition to the latent variable $\Theta$ and the observation $Z$:
@@
Here, $\{G_{\xi}\}_{\xi}$ is an unknown family of probability measures on $\mathbb{R}^{m}$ and $\{P_{\theta,\xi}\}_{\theta,\xi}$ is a known family of probability distributions on $\mathbb{R}^{d}$. We assume that $\Theta$ is unobserved while $(Z,\Xi)$ is observed. We also consider the EB setting where we have i.i.d. samples $(\Theta_{1},\Xi_{1},Z_{1}),\ldots,(\Theta_{n},\Xi_{n},Z_{n})$ from (5.1) and we observe $(\Xi_{1},Z_{1}),\ldots,(\Xi_{n},Z_{n})$.
In our application to astronomy in Subsection 6.1, we model measurements as $Z_{i}\sim\mathcal{N}(\Theta_{i},\Sigma_{i})$ where the covariance matrices $\Sigma_{i}$ vary across observations. This fits framework (5.1) by setting $\Xi_{i}=\Sigma_{i}$ and $P_{\theta,\xi}=\mathcal{N}(\theta,\xi)$, allowing us to handle the heterogeneous measurement uncertainty in astronomical data. In our baseball application in Subsection 6.2, part of our modeling involves Poisson observations with heterogeneous exposure, where we set $\Xi_{i}=\lambda_{i}$ and $P_{\theta,\xi}=\textnormal{Poi}(\xi\theta)$. We call $\lambda_{i}$ the exposure (Norberg, 1979), representing the number of games played by player $i$, so that $\Theta_{i}=\mathbb{E}[Z_{i}/\lambda_{i}]$ is the rate normalized by exposure time. In both applications, we assume that $\{G_{\xi}\}_{\xi}$ does not depend on $\xi$, i.e., there is some unknown $G$ satisfying $G_{\xi}=G$ for all $\xi$ [^1] Allowing for dependence of $G_{\xi}$ on $\xi$ is important in some applications, see Appendix D.1 and Chen (2022).
In Appendix D, we discuss applications of (5.1) in more detail (Appendix D.1) and describe modifications of the EB procedures in Section 4 to construct constrained denoisers under model (5.1). Constraints on the denoising procedure have two possible forms, namely a constraint on the marginal distribution of $\delta(Z,\Xi)$ or a constraint on the conditional distribution of $\delta(Z,\Xi)$ given $\{\Xi=\xi\}$ for almost every $\xi$. In particular, we discuss marginal and conditional variance constraints (Appendix D.2) leading to denoisers $\delta_{\mathcal{M}\mathcal{VCB}}$ and $\delta_{\mathcal{C}\mathcal{VCB}}$ respectively, marginal distribution constraints (Appendix D.3) leading to a denoiser $\delta_{\mathcal{M}\mathcal{DCB}}$, and marginal general constraints (Appendix D.4), leading to a denoiser $\delta_{\mathcal{M}\mathcal{GCB}}$; as usual, we write $\hat{\delta}_{\ast}$ to denote an EB approximation of an oracle denoiser $\delta_{\ast}$. See also Section D.2.3 for a numerical illustration of the difference between the marginal conditional variance constraints.